Solve for x: 4x^2-28=0
Answer by checkley77(12844)

4x^2-28=0 factor out the 4
4(x^2-7)=0
x^2-7=0 add 7
x^2=7 take the sqrt
x=sqrt7 or 2.64575 ans.